OpenStudy (anonymous):

HELP ? Explain how to find the solution to the system below. Then find the solution(s) to the system. Round answers to the nearest hundredth, if necessary. y=x^2-x-12 y=x

OpenStudy (anonymous):

Explanation: Because y = x as given by the second equation, you can substitute it into the first equation ===> you now have 1 equation with only x as variable ==> you can rearrange the terms to solve for x. Now can you implement that?

OpenStudy (anonymous):

how would u rearrange them ?

OpenStudy (anonymous):

hmm idk doesn't mean the x stays the same or you add them ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

ohh so it will also equal 1 ?

OpenStudy (mathstudent55):

\( y=x^2-x-12 \) \(y=x\) Substitute x in for y in the first equation: \( x=x^2-x-12 \) \(x^2 - 2x - 12 = 0 \) Use the quadractic formula, or complete the square to solve for x.

OpenStudy (anonymous):

sorry xoxox123 :(

OpenStudy (anonymous):

sorry for what ??? @mth3v4

OpenStudy (anonymous):

i have mislead you :P

OpenStudy (anonymous):

your good its ok.

OpenStudy (anonymous):

ax2 + bx + c = 0" @mathstudent55 so i substuite my numbers into this

OpenStudy (mathstudent55):

To use the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ax^2

OpenStudy (mathstudent55):

\( x^2 - 2x - 12 = 0\) \( ax^2 + bx + c = 0\) \( x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) Use \(a = 1\), \(b = -2\), and \(c = -12\) in the quadratic formula.

OpenStudy (anonymous):

-2+ √-2^2-(1)(-12)/2(1) ???????????

OpenStudy (mathstudent55):

Notice the first thing is -b which turns to -(-2) = 2 \(x = \dfrac{-(-2) \pm \sqrt{(-2)^2 - 4(1)(-12)}}{2(1)}\)

OpenStudy (anonymous):

so what next after u do that

OpenStudy (mathstudent55):

Simplify wherever you can one step at a time until you reach a value for x.

OpenStudy (mathstudent55):

\(x = \dfrac{-(-2) \pm \sqrt{(-2)^2 - 4(1)(-12)}}{2(1)}\) \(x = \dfrac{2 \pm \sqrt{4 - (-48)}}{2}\) \(x = \dfrac{2 \pm \sqrt{4 + 48}}{2}\) \(x = \dfrac{2 \pm \sqrt{52}}{2}\) \(x = \dfrac{2 \pm \sqrt{4 \times 13}}{2}\) \(x = \dfrac{2 \pm 2\sqrt{13}}{2}\) \(x = 1 \pm \sqrt{13}\)

OpenStudy (anonymous):

cant u simplify any further than that ?

OpenStudy (mathstudent55):

Since the problem wants the answers rounded off, you can do this: \(x = 1 + \sqrt{13} = 4.61 \) or \(x = 1 - \sqrt{13} = -2.61 \) Remember that this is a system of equations, so you need the solutions as ordered pairs. Since the second equation is y = x, for each value of x, we have a value of y that is equal to it: The solutions are: (4.61, 4.61) and (-2.61, -2.61)